# Online Iterative Reinforcement Learning from Human Feedback with General Preference Model

### v2

@@ 6.1 Practical Algorithmic Design
where $\mathrm{SG}[\pi]$ means that although we generate data from policy $\pi$, but we do not compute the gradient for this data-generation process. We have the following result.
Proposition 1 (Proposition 4.1 of Calandriello et al., (2024)).
The minimizer of Equation (12) is the unique Nash policy of the Equation (9).
Enhancer explores via rejection sampling. According to Equation (11), the enhancer aims to find a policy that 1) is close to the main agent’s policy $\hat{\pi}_{t}^{1}$ but also 2) maximizes the uncertainty relative to the $\hat{\pi}_{t}^{1}$. However, since for the general neural network, the uncertainty estimator does not admit a closed form, in practice, we typically resort to heuristic methods. One popular way to do so in the context of alignment is the rejection sampling (Nakano et al.,, 2021; Dong et al.,, 2023; Liu et al., 2023a, ; Hoang Tran,, 2024; Yuan et al.,, 2024). Specifically, given a prompt $x$, we use $\hat{\pi}_{t}^{1}$ to independently sample $n$ responses, use a tournament-style procedure to get the best response (and reject all other responses), and take the best responses as $\hat{\pi}_{t}^{2}$. In other words, we take the policy-induced by rejection sampling with $\hat{\pi}_{t}^{1}$ and $P^{*}$ as the enhancer policy $\hat{\pi}_{t}^{2}$. In this way, the $\hat{\pi}_{t}^{2}$ enlarges the margins between $\hat{\pi}_{t}^{1}$ while maintaining a moderate KL divergence. For instance, in the special case of the BT model, if we rank the samples via the learned reward, the KL divergence is upper bounded by $\log n-\frac{n-1}{n}$ and is usually far better than this conservative estimation (Beirami et al.,, 2024).
6.2 Preference Model Construction
Bradley-Terry model construction. We follow the previous works (Ouyang et al.,, 2022; Bai et al., 2022a, ) to initialize the reward model using an SFT model but replace the last layer with a linear head to predict a scalar score. The loss function of reward modeling is the negative log-likelihood so that minimizing the loss is equivalent to MLE:
where $a^{w}$ is the preferred response over $a^{l}$. We use the open-source training script [^4] https://github.com/WeiXiongUST/RLHF-Reward-Modeling. We train the model for one epoch and use a batch size of 256, a learning rate of $\mathrm{lr}=$ 1e-5, and a cosine learning rate schedule with a warm-up ratio of 0.03.
Preference model construction. We follow Zhao et al., (2023); Liu et al., 2023a to utilize the fact that the LLM is the next token predictor for the preference modeling. Specifically, suppose that we have a preference pair $(x,a^{1},a^{2},A)$ where $A$ means that the first response is better. Then, we format it as
@@ 6.2 Preference Model Construction
Then, we simply treat the preference modeling as an instruction-following task to fine-tune the model on these instruction-label pairs. In particular, to mitigate the position bias (the preference model may prefer the response that is given in the position of RESPONSE A), we randomly switch the order of the two responses in the data formatting process. During inference, we simply use the probability of decoding A as the $\hat{P}(x,a^{1},a^{2})$. We mention in passing that it is also possible to include a rubric in the instruction template to guide the model’s prediction and achieve better results (Qin et al.,, 2023). We do not explore the prompt engineering for the best preference model construction for simplicity.
Ground-truth preference model for simulation. Ideally, the $P^{*}$ is supposed to be a group of human labelers or closed-source LLMs like Chat-GPT. Unfortunately, due to resource constraints, we cannot afford the cost of using these preference oracles. Instead, we follow Gao et al., (2023) to use a strong preference model to serve as the $P^{*}$ in the simulation. Specifically, we adopt the LLaMA3-8B, and train the preference model on a diverse set of open-source preference datasets including HH-RLHF (Bai et al., 2022a, ), Stanford Human Preferences Dataset (SHP) (Ethayarajh et al.,, 2022), Ultra-feedback (Cui et al.,, 2023), HelpSteer (Wang et al., 2023c, ), distilabel-capybara [^5] https://huggingface.co/datasets/argilla/distilabel-capybara-dpo-7k-binarized, distilabel-orca [^6] https://huggingface.co/datasets/argilla/distilabel-intel-orca-dpo-pairs, and UltraInteract777openbmb/UltraInteract_pair. Motivated by the Theorem 1 as well as the practical application (Ouyang et al.,, 2022), we include more than 1 comparison pair when a prompt is with more than 2 responses for better coverage. To be specific,
  • for UltraInteract, we take a subset of 150K pairs into the mixture.
We have about 700K preference pairs in the final training set and we have uploaded it to the huggingface at https://huggingface.co/datasets/hendrydong/preference_700K. We use the package [^8] https://github.com/OpenAccess-AI-Collective/axolotl to perform supervised finetuning, with the detailed hyper-parameters given in Appendix E. The resulting preference models are evaluated by the reward bench (Lambert et al.,, 2024), with the results summarized in Table 1. The preference model based on LLaMA3-8B-it achieves state-of-the-art test accuracy and can serve as a stable preference oracle for the simulation study.
@@ 7 Conclusion
8 Acknowledgement
The authors would like to thank Tianqi Liu for insightful discussions on the training of preference model, and thank Haoxiang Wang, and Zihao Li for valuable discussions on the preference dataset selection.
Appendix A Notation Table
@@
$J(\pi)$ The KL-regularized target defined in Equation (3).
$\eta$ The coefficient of KL penalty, defined in Equation (3).
$\ell_{\mathcal{D}}$ The log-likelihood function defined in Equation (6).
@@
$\mathcal{C}(\pi,\pi_{D},\mathcal{P})$ Coverage term for version space based Algorithm 1 defined in Theorem 1.
@@
$\Gamma(\lambda,\pi^{1},\pi^{2})$ Infomation ratio defined in Definition Equation (3).
$\widetilde{\Gamma}_{t}^{m}(\lambda,\pi^{1},\pi^{2}),\Gamma(x,\pi^{1},\pi^{2})$ Uncertainty bonus defined in Equation (10) and Equation (18).
@@
Table 2: The table of notations used in this paper.
Appendix B Proofs for the Offline Setting
B.1 Proof for Theorem 1
Lemma 2 (Restatement of Lemma 1).
@@ B.1 Proof for Theorem 1
We continue to lower bound the Hellinger by
@@
where the inequality uses the fact that for any distribution $p,q$, $H(p,q)\geq\mathrm{TV}(p,q)$ according to Theorem B.9 of Zhang, (2023).
@@
where the second inequality uses Equation (B.1), and the last inequality uses Equation (B.1). By taking $P$ as $\hat{P}$, since $\hat{P}$ is the MLE, we get
@@
The following decomposition holds
$J(\pi_{*}^{1},\pi^{2}_{*})-J(\hat{\pi}^{1},\dagger)\leq$ $\underbrace{\underline{J}(\hat{\pi}^{1},\widetilde{\pi}^{2})-J(\hat{\pi}^{1},\dagger)}_{\displaystyle q_{1}}+\underbrace{\underline{J}(\pi_{*}^{1},\widetilde{\pi}^{2})-\underline{J}(\hat{\pi}^{1},\widetilde{\pi}^{2})}_{\displaystyle q_{2}}$
$+$ $\underbrace{J(\pi_{*}^{1},\widetilde{\pi}^{2})-\underline{J}(\pi_{*}^{1},\widetilde{\pi}^{2})}_{\displaystyle q_{3}}+\underbrace{J(\pi_{*}^{1},\pi_{*}^{2})-J(\pi_{*}^{1},\widetilde{\pi}^{2})}_{\displaystyle q_{4}}.$
Then, we bound these terms separately. For the term $q_{1}$,
$q_{1}=$ $\underline{J}(\hat{\pi}^{1},\widetilde{\pi}^{2})-\min_{\pi_{2}}\mathbb{E}_{x\sim d_{0}}\big{[}P^{*}(x,\hat{\pi}^{1},\pi^{2})-\eta^{-1}D_{\mathrm{KL}}(\hat{\pi}^{1}(\cdot|x)\|\pi_{0}(\cdot|x))+\eta^{-1}D_{\mathrm{KL}}(\pi^{2}(\cdot|x)\|\pi_{0}(\cdot|x))\big{]}$
$\leq$ $\underline{J}(\hat{\pi}^{1},\widetilde{\pi}^{2})-\min_{\pi_{2}}\min_{P\in\mathcal{P}}\mathbb{E}_{x\sim d_{0}}\big{[}P^{*}(x,\hat{\pi}^{1},\pi^{2})-\eta^{-1}D_{\mathrm{KL}}(\hat{\pi}^{1}(\cdot|x)\|\pi_{0}(\cdot|x))+\eta^{-1}D_{\mathrm{KL}}(\pi^{2}(\cdot|x)\|\pi_{0}(\cdot|x))\big{]}$
$=$ $0$
For the term $q_{2}$, since $(\hat{\pi}^{1},\widetilde{\pi}^{2})$ is the Nash equilibrium of $\underline{J}$, we have
$q_{2}=\underline{J}(\pi_{*}^{1},\widetilde{\pi}^{2})-\underline{J}(\hat{\pi}^{1},\widetilde{\pi}^{2})\leq 0.$
For the term $q_{3}$,
$q_{3}=$ $\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a^{1}\sim\pi_{*}^{1},a^{2}\sim\widetilde{\pi}^{2}}P^{*}(x,a^{1},a^{2})-\min_{P\in\widehat{\mathcal{P}}}\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a^{1}\sim\pi_{*}^{1},a^{2}\sim\widetilde{\pi}^{2}}P(x,a^{1},a^{2})$
$=$ $\min_{P\in\widehat{\mathcal{P}}}\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a^{1}\sim\pi_{*}^{1},a^{2}\sim\widetilde{\pi}^{2}}[\hat{P}(x,a^{1},a^{2})-P(x,a^{1},a^{2})]+\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a^{1}\sim\pi_{*}^{1},a^{2}\sim\hat{\pi}^{2}}[P^{*}(x,a^{1},a^{2})-\hat{P}(x,a^{1},a^{2})]$
$\leq$ $2\beta\widetilde{\Gamma}(\pi_{*}^{1},\widetilde{\pi}^{2}),$
@@
$\widetilde{\Gamma}(\pi^{1},\pi^{2}):=\sup_{P\in\mathcal{P}}\frac{|\mathbb{E}_{x\sim d_{0}}[P(x,\pi^{1},\pi^{2})-\hat{P}(x,\pi^{1},\pi^{2})]|}{\sqrt{\lambda+\sum_{i=1}^{n}(P(x_{i},a_{i}^{1},a_{i}^{2})-\hat{P}(x_{i},a_{i}^{1},a_{i}^{2}))^{2}}}.$
Since $\pi_{*}^{2}$ is the best response to $\pi_{*}^{1}$ with respect to $J$, we have $q_{4}\leq 0$. Therefore, we obtain that
$J(\pi_{*}^{1},\pi^{2}_{*})-J(\hat{\pi}^{1},\dagger)\leq 2\beta\widetilde{\Gamma}(\pi_{*}^{1},\widetilde{\pi}^{2}).$ (15)

### v3

@@ 5 Practical Implementation of Preference Model and Iterative RLHF
where $\mathrm{SG}[\pi]$ means that although we generate data from policy $\pi$, but we do not compute the gradient for this data-generation process. Moreover, according to Proposition 4.1 of Calandriello et al. [11], the minimizer of (13) is the unique Nash policy of the (10).
Enhancer explores via rejection sampling. According to (12), the enhancer aims to find a policy that (1) is close to the main agent’s policy $\hat{\pi}_{t}^{1}$; (2) maximizes the uncertainty relative to the $\hat{\pi}_{t}^{1}$. However, since for the general neural network, the uncertainty estimator does not admit a closed form, in practice, we typically resort to heuristic methods. One popular way in the context of alignment is the rejection sampling [47, 18, 43, 31, 76]. Specifically, given a prompt $x$, we use $\hat{\pi}_{t}^{1}$ to independently sample $n$ responses, use a tournament-style procedure to get the best response (and reject all other responses), and take the best responses as $\hat{\pi}_{t}^{2}$. In other words, we take the policy induced by rejection sampling with $\hat{\pi}_{t}^{1}$ and $P^{*}$ as the enhancer policy $\hat{\pi}_{t}^{2}$. In this way, the $\hat{\pi}_{t}^{2}$ enlarges the margins between $\hat{\pi}_{t}^{1}$ while maintaining a moderate KL divergence. For instance, in the special case of the BT model, if we rank the samples via the learned reward, the KL divergence is upper bounded by $\log n-(n-1)/n$ and is usually far better than this conservative estimation [6].
Preference model construction. We follow Zhao et al. [85], Liu et al. [43], Dong et al. [19] to utilize the fact that the LLM is the next token predictor for the preference modeling. Specifically, we have a preference pair $(x,a^{1},a^{2},A)$, where $A$ means that the first response is better, which is formatted as
@@
Then, we simply treat the preference modeling as an instruction-following task to fine-tune the model on these instruction-label pairs. In particular, to mitigate the position bias (the preference model may prefer the response that is given in the position of RESPONSE A), we randomly switch the order of the two responses in the data formatting process. During inference, we simply use the probability of decoding A as the $\hat{P}(x,a^{1},a^{2})$. We mention in passing that it is also possible to include a rubric in the instruction template to guide the model’s prediction and achieve better results [52]. We observe the benefits of the additional prompt engineering in early experiments but decide to use the current version because the main focus is to verify the effectiveness of general preference structure. This implementation is also referred to as the Generative RM in subsequent works.
6 Experiments
Model, Dataset, and Evaluation. We adopt the widely used open-source model Zephyr-SFT-7B [61] as the starting checkpoint, which is based on the Mistral-7B-v0.1444https://huggingface.co/mistralai/Mistral-7B-Instruct-v0.1 and fine-tuned on 200K high-quality Ultra-chat data [16]. We use the Ultra-feedback [16] as our prompt set. We divide the prompt set into the train set (60K), validation set (1K), and test set (3K). We mainly use head-to-head comparisons to evaluate the resulting models. In particular, we consider two types of win rate: 1) the win rate measured by the ground-truth LLaMA3-8B-based preference model on the hand-out test set from UltraFeedback; 2) the win rate measured by the GPT-4 Preview (11/06) on an out-of-distribution prompt set AlpacaEval2 [21]. Specifically, for the first evaluation, we use the best DPO model as the reference model, and for the AlpacaEval2, the GPT-4 Preview (11/06) is used as a reference model, and as the judge at the same time.
Method and Competitors. We consider the implementation of Algorithm 2 with self-play IPO and rejection sampling as discussed in Section 5. We iterate for three iterations in total and for each iteration, we retrain a preference model using all the historical data, and run self-play IPO from the initial checkpoint $\pi_{0}$ (i.e., Zephyr-7B-SFT). For simplicity, we refer to this algorithm as Online ELHF IPO. We use the offline DPO [53], offline IPO [3], and SFT model as the baseline. In particular, we do not further fine-tune the Zephyr-7B-SFT on the preferred responses of Ultra-Feedback because the quality of Ultra-Feedback is lower than that of Ultra-Chat, which is generated by Chat-GPT APIs. For DPO, we follow Xiong et al. [72], Tunstall et al. [61], Rafailov et al. [53] to set the KL coefficient as $\eta=0.1$. For IPO, we search the hyper-parameter in $\{0.1,0.5,1.0\}$ and report the results in Table 2. Clearly, the model with $\eta=0.1$ beats all other IPO models and the SFT model with large margins, so we set $\eta=0.1$ for the offline IPO and the Online ELHF IPO algorithm in the subsequent studies.
Table 2: The evaluation results of the IPO-aligned models under different KL coefficients. For the first 4 win rates, we use the LLaMA3-8B-based preference model to conduct head-to-head comparisons on the hand-out test set from Ultra-feedback with 3K prompts.
Models v.s. SFT v.s. $\eta=0.1$ v.s. $\eta=0.5$ v.s. $\eta=1.0$ AlpacaEval2
SFT $0.5$ 0.121 0.205 0.231 4.63
Offline-IPO-$\eta=0.1$ 0.879 0.5 0.673 0.769 9.36
Offline-IPO-$\eta=0.5$ 0.795 0.327 0.5 0.632 6.86
Offline-IPO-$\eta=1.0$ 0.710 0.230 0.328 0.5 6.55
Table 3: The evaluation results of the models from different RLHF algorithms. The gold win rates are computed on the hand-out test set from Ultra-feedback with 3K prompts, with the Offline DPO model as the reference. Details of AlpacaEval2 can be found in Dubois et al. [21].
Models Settings Gold WR v.s. IPO AlpacaEval2 WR
SFT - 0.121 4.63
Offline DPO Offline 0.41 9.33
Offline IPO Offline 0.5 9.36
Online-ELHF-IPO Online 0.78 17.67
Simulation framework. For all the offline algorithms, we sample two responses per prompt of the train set and use the LLaMA3-8B-based preference model to give the preference signal. Then, we run offline DPO and IPO with the synthetic dataset. For the Online ELHF, we set $n=4$ in the rejection sampling process and use a tournament-style ranking method (so that the complexity of rejection sampling is linear in $n$) to find the best response.
IPO, DPO, and Online ELHF-IPO. We use the open-source project TRL [^5] https://github.com/huggingface/trl to implement IPO and DPO. In particular, we have implemented IPO with log-likelihood/perplexity (perplexity is averaged log-likelihood by sequence length), where the original authors of IPO suggest that log-likelihood-based implementation is unstable (see the huggingface blog [^6] https://huggingface.co/blog/pref-tuning for details). We also found that the IPO without average cannot normally converge and is of poor performance and take the perplexity implementation accordingly. For DPO, we implement the vanilla version as the baseline. We present the main result in Table 3. It is clear that Online ELHF-IPO outperforms the baselines.
7 Related Work
This section focuses on the theoretical aspects. A general discussion is provided in Appendix B.1
Theoretical Study of Reward-based RLHF. The theoretical study of policy optimization from preference feedback dated back to the dueling bandits [e.g., 78, 55, 7]. This was later extended to the online RL setting by Xu et al. [73], Novoseller et al. [48], Pacchiano et al. [51], Chen et al. [12], including tabular online RLHF with finite state, and general function approximation for capturing real problems with large state spaces. Zhan et al. [81], Wu and Sun [68] further encompasses the development of reward-free learning type algorithms and sampling-based algorithms for online RLHF. Apart from the online setting, there is another line of works [88, 80, 40] studying the reward-based RLHF in the offline setting, which learns from a pre-determined offline dataset with suitable coverage condition over the state-action space. However, they consider reward maximization and deviate from the practical applications (e.g., these frameworks admit a deterministic optimal policy). Recently, Xiong et al. [72] first formulated the RLHF as a reverse-KL regularized contextual bandit and provided finite-sample guarantees in offline, online, and hybrid settings. We remark that all these papers consider only the reward-based RLHF framework, thus differing from ours.
Theoretical Study of RLHF under General Preference Oracle. Our work is related to Dudík et al. [22] and Wang et al. [65]. They investigate preference-based RLHF under a general preference model. The major difference is that we consider the reverse-KL regularized preference, aligning closely with recent LLM advancements [90, 50, 4, 53], while previous work only considers the non-regularized one. Meanwhile, Dudík et al. [22] considers the problem of finite action, while our work and Wang et al. [65] consider the problem with large or even infinite state-action under function approximation. In terms of learning paradigm and algorithmic design, we consider both offline learning from a pre-collected dataset and batch online learning with a sparse policy update, while Dudík et al. [22], Wang et al. [65] studies sequential online learning that updates policy in each step, which is not feasible in the context of LLMs. Moreover, we demonstrate that the proposed algorithms can be reasonably implemented in practice, but Dudík et al. [22], Wang et al. [65] only focus on information-theoretical algorithms. To summarize, the framework in this work accurately reflects real-world alignment practices thus aligning more closely with the RLHF practice. Our work is closely related to the IPO [3] and Nash learning [46], which also motivate new algorithmic design with a general preference oracle. We comment on the similarities and differences between our framework and theirs as follows. In terms of the problem setting, our work and Nash learning consider the minimax game under the reverse-KL regularized preference, while IPO can be interpreted to find the best response of the fixed reference policy, and may be considered as a special case of the game formulation. In terms of learning paradigm, both the IPO and Nash learning only consider learning toward a fixed and known preference oracle, and study the optimization property of the problem: how to compute the optimal policy under the given preference oracle. In contrast, we study the statistical property, where the preference model needed to be learned and our goal is to find the optimal policy under the underlying ground-truth preference model. In particular, the computational challenge is hidden in Definition 2 and Munos et al. [46] provides a reasonable approximation of the planning oracle. In this sense, our work and Munos et al. [46] are complementary to each other. Finally, the concurrent work Swamy et al. [58] studies the non-regularized general preference model in the sequential online setting and aims to find the Nash equilibrium in the context of continuous control tasks. In terms of the observation model, they assume access to the preference score $\mathbb{P}(a^{1}\succ a^{2}|x,a^{1},a^{2})$, while we only observe the preference signal $y\sim\mathrm{Ber}(\mathbb{P}(a^{1}\succ a^{2}|x,a^{1},a^{2}))$. Moreover, they design online RLHF algorithms based on a reduction to the no-regret algorithm like Hedge [25], whose techniques are fundamentally different from ours.
@@ 8 Conclusion
9 Acknowledgment
The authors would like to thank Tianqi Liu for insightful discussions on the training of the preference model, and thank Haoxiang Wang, and Zihao Li for valuable discussions on the preference dataset selection. We also thank Nevena Lazic and Csaba Szepesvari for pointing out a technical gap in the first version.
Wei Xiong and Tong Zhang are partially supported by an NSF IIS grant No. 2416897 and Tong Zhang is partially supported by an NSF IIS grant No. 2416897. Nan Jiang acknowledges funding support from NSF IIS-2112471, NSF CAREER IIS-2141781, Google Scholar Award, and Sloan Fellowship.
Appendix A Authorship and Credit Attribution
All authors provided valuable contributions to this project, each bringing unique expertise and insights that were crucial for its success.
  • CY investigated the general preference problem, proved the theoretical results for both offline and online settings, and wrote the main part of the paper.
  • WX first proved the effectiveness of the general preference model, proposed the online iterative algorithm, contributed to the proof and paper writing, and contributed to the experiment.
  • YZ proved the properties of the general preference problem, made important contributions to the offline result and the proof, contributed to the paper writing.
  • HD designed the practical implementation under generalized preference model and conducted most experiments to show the effectiveness of the proposed algorithm.
  • NJ and TZ supported and advised the junior authors’ works, provided computational resources and suggested experiments and writings.
Appendix B Notation Table, Related Work, Experimental Details
Table 4: The table of notations used in this paper.
@@ Appendix B Notation Table, Related Work, Experimental Details
$J(\pi)$ The KL-regularized target defined in (3).
$\eta$ The coefficient of KL penalty, defined in (3).
$\ell_{\mathcal{D}}$ The log-likelihood function defined in (7).
@@
$\mathcal{C}(\pi,\pi_{D},\mathcal{P})$ Coverage term for version space-based Algorithm 1 defined in Theorem 1.
@@
$\Gamma(\lambda,\pi^{1},\pi^{2})$ Information ratio defined in Definition (3).
$\widetilde{\Gamma}_{t}^{m}(\lambda,\pi^{1},\pi^{2}),\Gamma(x,\pi^{1},\pi^{2})$ Uncertainty bonus defined in (11) and (19).
@@
B.1 More Related Work
RLHF. RLHF was first popularized in the deep RL literature by Christiano et al. [14], which served to direct the attention of the RL community to the preference-based feedback, but may further date back to Bennett et al. [8], Knox and Stone [36] in the context of machine learning. It has attracted significant attention recently, mainly due to its tremendous success in Chat-GPT [49]. The most popular and standard RLHF framework is outlined in Ouyang et al. [50], Touvron et al. [60] and we have described the details in Section 1. In terms of reward optimization, PPO [56] is the most well-known algorithm in LLM alignment literature. However, tuning the PPO algorithm to the best performance requires extensive efforts and the result of Chat-GPT4 [49] has not been widely reproduced so far. This motivates another line of works of algorithms that are based on supervised learning. For instance, Dong et al. [18], Yuan et al. [77], Touvron et al. [60], Gulcehre et al. [30], Ji et al. [33] propose reward ranked finetuning, (also known as rejection sampling finetuning), which essentially learns from the best-of-n policy [47] to maximize the reward. The reward-ranked finetuning algorithm is a stable policy optimization algorithm with minimal hyper-parameter configuration and was applied to the RLHF of LLaMA2 [60]. However, it is also observed that the reward ranked finetuning algorithm leads to considerable forgetting in a wide range of tasks (also referred to as the alignment tax), as the algorithmic design only considers reward optimization [60, 42, 13]. One approach to mitigate this issue is to use the KL-regularized formulation, which is widely adopted in the deep RL approach (e.g. PPO) [90, 67, 50, 4, 37, 41], and other supervised-learning-based algorithms [53, 63, 43, 3], whose theoretical property is studied in Xiong et al. [72]. Among them, (offline) Direct Preference Optimization (DPO) [53] has emerged as an attractive alternative approach to PPO with notable stability and competitive performance. Xiong et al. [72], Hoang Tran [31], Yuan et al. [76] further extend the offline DPO to the iterative (online) variant, and the resulting models demonstrate impressive performance [31, 19]. However, all these algorithms are designed under the reward-based RLHF framework to maximize the underlying reward function (with appropriate regularization).
B.2 Details of Experiments
Bradley-Terry model construction. We follow the previous works [50, 4] to initialize the reward model using an SFT model but replace the last layer with a linear head to predict a scalar score. The loss function of reward modeling is the negative log-likelihood so that minimizing the loss is equivalent to MLE:
where $a^{w}$ is the preferred response over $a^{l}$. We train the model for one epoch and use a batch size of 256, a learning rate of $\mathrm{lr}=$ 1e-5, and a cosine learning rate schedule with a warm-up ratio of 0.03.
Ground-truth preference model for simulation. Ideally, the $P^{*}$ is supposed to be a group of human labelers or closed-source LLMs like Chat-GPT. Unfortunately, due to resource constraints, we cannot afford the cost of using these preference oracles. Instead, we follow Gao et al. [26] to use a strong preference model to serve as the $P^{*}$ in the simulation. Specifically, we adopt the LLaMA3-8B, and train the preference model on a diverse set of open-source preference datasets including HH-RLHF [4], Stanford Human Preferences Dataset (SHP) [23], Ultra-feedback [16], HelpSteer [66], distilabel-capybara [^7] https://huggingface.co/datasets/argilla/distilabel-capybara-dpo-7k-binarized, distilabel-orca [^8] https://huggingface.co/datasets/argilla/distilabel-intel-orca-dpo-pairs, and UltraInteract999openbmb/UltraInteract_pair. Motivated by the Theorem 1 as well as the practical application [50], we include more than 1 comparison pair when a prompt is with more than 2 responses for better coverage. To be specific,
  • for UltraInteract, we take a subset of 150K pairs into the mixture.
We have about 700K preference pairs in our training stage. We use the package axolotl [^10] https://github.com/OpenAccess-AI-Collective/axolotl to perform supervised fine-tuning, with the detailed hyper-parameters given in Appendix B.2. The resulting preference models are evaluated by the reward bench [39], with the results summarized in Table 1. The preference model based on LLaMA3-8B-it achieves state-of-the-art test accuracy and can serve as a stable preference oracle for the simulation study.
We present the hyper-parameters in Table 5. All experiments are conducted on 8$\times$×A100-40G with Deepspeed ZeRO-3.
Models Hyper-parameter Value
Learning rate $1\times 10^{-5}$
Scheduler Cosine decay with 0.03 warm-up
Reward model with Gemma-2B-it Epoch 1
Batch size 256
Learning rate $1\times 10^{-5}$
Scheduler Cosine decay with 0.03 warm-up
Epoch 1
Preference model with Gemma-2B-it Batch size 128
Packing True
Block size 3072
Learning rate $1\times 10^{-5}$
Scheduler Cosine decay with 0.03 warm-up
Epoch 1
Preference model with LLaMA3-8B-it Batch size 128
Packing True
Block size 3072
Examples from Ultra-feedback.
We provide several examples here:
  • Create a list of three mistakes to avoid when designing an AI assistant.
  • Pretend you’re a next.js expert, write ad copy about a free trial on Vercel.
  • Can you describe the role of photography in shaping the art world?
Appendix C Proofs for the Offline Setting
C.1 Proof for Theorem 1
Lemma 1.
@@ C.1 Proof for Theorem 1
We continue to lower-bound the Hellinger by
@@
where the inequality uses the fact that for any distribution $p,q$, $H(p,q)\geq\mathrm{TV}(p,q)$ according to Theorem B.9 of Zhang [83].
@@
where the second inequality uses (C.1), and the last inequality uses (C.1). By taking $P$ as $\hat{P}$, since $\hat{P}$ is the MLE, we get
@@
Let $\widetilde{\pi}^{2}_{*}=\min_{\pi^{2}\in\Pi}\underline{J}(\pi^{1}_{*},\pi^{2})$ and $\pi^{\dagger,2}=\min_{\pi^{2}\in\Pi}J(\hat{\pi}^{1},\pi^{2})$. The following decomposition holds
$J(\pi_{*}^{1},\pi^{2}_{*})-J(\hat{\pi}^{1},\pi^{\dagger,2})\leq$ $\underbrace{J(\pi_{*}^{1},\pi_{*}^{2})-J(\pi_{*}^{1},\widetilde{\pi}^{2}_{*})}_{\displaystyle q_{1}}+\underbrace{J(\pi_{*}^{1},\widetilde{\pi}^{2}_{*})-\underline{J}(\pi_{*}^{1},\widetilde{\pi}^{2}_{*})}_{\displaystyle q_{2}}+\underbrace{\underline{J}(\pi_{*}^{1},\widetilde{\pi}^{2}_{*})-\underline{J}(\hat{\pi}^{1},\widetilde{\pi}^{2})}_{\displaystyle q_{3}}$
$+$ $\underbrace{\underline{J}(\hat{\pi}^{1},\widetilde{\pi}^{2})-\underline{J}(\hat{\pi}^{1},\pi^{\dagger,2})}_{\displaystyle q_{4}}+\underbrace{\underline{J}(\hat{\pi}^{1},\pi^{\dagger,2})-J(\hat{\pi}^{1},\pi^{\dagger,2})}_{\displaystyle q_{5}}.$
Then, we bound these terms separately. For the term $q_{1}$, since $(\pi_{*}^{1},\pi_{*}^{2})$ is the Nash equilibrium of $J$, we have $q_{1}\leq 0$. For the term $q_{2}$,
$q_{2}=$ $\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a^{1}\sim\pi_{*}^{1},a^{2}\sim\widetilde{\pi}_{*}^{2}}P^{*}(x,a^{1},a^{2})-\min_{P\in\widehat{\mathcal{P}}}\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a^{1}\sim\pi_{*}^{1},a^{2}\sim\widetilde{\pi}_{*}^{2}}P(x,a^{1},a^{2})$
$=$ $\min_{P\in\widehat{\mathcal{P}}}\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a^{1}\sim\pi_{*}^{1},a^{2}\sim\widetilde{\pi}_{*}^{2}}[\hat{P}(x,a^{1},a^{2})-P(x,a^{1},a^{2})]+\mathbb{E}_{x\sim d_{0}}\mathbb{E}_{a^{1}\sim\pi_{*}^{1},a^{2}\sim\hat{\pi}_{*}^{2}}[P^{*}(x,a^{1},a^{2})-\hat{P}(x,a^{1},a^{2})]$
$\leq$ $2\beta\widetilde{\Gamma}(\pi_{*}^{1},\widetilde{\pi}_{*}^{2}),$
@@
$\widetilde{\Gamma}(\pi^{1},\pi^{2}):=\sup_{P\in\widehat{\mathcal{P}}}\frac{|\mathbb{E}_{x\sim d_{0}}[P(x,\pi^{1},\pi^{2})-\hat{P}(x,\pi^{1},\pi^{2})]|}{\sqrt{\lambda+\sum_{i=1}^{n}(P(x_{i},a_{i}^{1},a_{i}^{2})-\hat{P}(x_{i},a_{i}^{1},a_{i}^{2}))^{2}}}.$
By the optimality of $\hat{\pi}^{1}$, term $q_{3}\leq 0$. Since $\widetilde{\pi}^{2}$ is the best response to $\hat{\pi}^{1}$ with respect to $\underline{J}$, we have $q_{4}\leq 0$. From Lemma 1, we know that $P^{\star}\in\widehat{\mathcal{P}}$, thus $q_{5}\leq 0$. Putting everything together, we obtain that
$J(\pi_{*}^{1},\pi^{2}_{*})-J(\hat{\pi}^{1},\pi^{\dagger,2})\leq 2\beta\widetilde{\Gamma}(\pi_{*}^{1},\widetilde{\pi}_{*}^{2}).$ (16)